How many molecules of zinc sulfate monohydrate does a 5mL of vitamin have?


1
Expert's answer
2021-01-11T03:46:03-0500

n = m/M

m (ZnSO4.H2O) = 27.44/1000 = 0.02744 g

M (ZnSO4.H2O) = 179.49 g/mol

n (ZnSO.H2O) = 0.02744/179.49 = 0.00015 mol


Each mole of matter contains 6.02214076×1023 molecules.


Given amount of ZnSO.H2O is equal to: 0.00015 x 6.02214076×1023 = 1020 molecules.


1020 molecules of zinc sulfate monohydrate does a 5mL of vitamin have